ideal airways is a small local carrier that flies among the northeast states. The company provides only one class of seats: coach. The following data are available:

Average full passenger fare- $150

Number of seats per plane- 120

Average load factor- 70%

Average variable cost per passenger- $40

Fixed cost per month- $1,800,000.

A. Determine the break even point in passengers and in revenues.

B. Determine the breakeven point in number of flights

C. If ideal raises its average full passenger fare to $200, it is estimated that the load factor will decrease to 55%. in this case what will be the breakeven point in number of flights

D. The cost of fuel is a significant variable cost to any airline. if fuel charges increase $8 per barrel, it is estimated that variable cost per passenger will rise to $60. in this case, what will be the new breakeven point in passengers and in number of flights.

E. Ideal has experienced an increase in variable cost per passenger to $50 and an increase in total fixed operating cost of 2,000,000. The company has decided to raise the average fare to $180. What number of passengers is needed to generate a pre-tax incomes of $1,000,000.

F. Ideal is considering offering a discounted fare of $120, which the company feels would increase the load factor to 80%. Only the additional seats would be sold at the discounted fare. Additional monthly advertising costs would be $100,000. How much before-tax income would the discounted fare provide ideal airways if the company has 20 flights per day, 30 days per month?
Answer #1

As per policy, only four parts of a question is allowed to answer at a time, but answering 5 parts here for you thanks:

A) Break Even Point (in units) = Fixed Cost / (Revenue per unit - VC per unit)
Break Even Point (in units) = 1800000 / (150 - 40)= 16364 passengers
Break Even Point (in revenue) = 1800000 *150 / (150 - 40)= 2454600
B) Break Even Point (in number of flights) = BEP (in passengers) / number of seats per plane * load factor
Break Even Point (in number of flights) =16364 / (120* 70%) = 194.81 or 195 flights
C) Break Even Point (in units) = 1800000 / (200 - 40)= 11250 passengers
Break Even Point (in number of flights) =11250 / (120* 55%) = 170.45 or 171 flights
D) Break Even Point (in units) = 1800000 / (150 - 60)= 20000 passengers
Break Even Point (in number of flights) =20000 / (120* 70%) = 238.1 or 238 flights
E) Break Even Point (in units) = (Fixed Cost + Target profit)/ (Revenue per unit - VC per unit)
Break Even Point (in units) = (2000000 + 1000000)/ (180 - 50)= 23077 passengers